# Large deviations in statistics of the local time and occupation time for a run and tumble particle

### v1

@@ Appendix B Calculation of SCGF for the occupation time statistics
$k-\lambda(k)=\gamma-\sqrt{\gamma^{2}-B}$ (105)
@@
$\lambda=\frac{2\gamma\left(\gamma-\sqrt{\gamma^{2}-B}\right)\sin^{2}\left(\frac{\sqrt{B}a}{v_{0}}\right)}{\sqrt{\gamma^{2}-B}+\gamma\cos\left(\frac{2\sqrt{B}a}{v_{0}}\right)}$ (107)
@@
$k-\lambda(k)=\gamma+\sqrt{\gamma^{2}-B}$ (108)
@@
$\lambda=\frac{2\gamma\left(\sqrt{\gamma^{2}-B}+\gamma\right)\sin^{2}\left(\frac{\sqrt{B}a}{v_{0}}\right)}{\gamma\cos\left(\frac{2\sqrt{B}a}{v_{0}}\right)-\sqrt{\gamma^{2}-B}}\,.$ (109)
@@
$k-\lambda(k)=\gamma+\sqrt{\gamma^{2}+|B|}$ (111)
@@
$\lambda=\frac{2\,\gamma\left(\sqrt{|B|+\gamma^{2}}+\gamma\right)\sinh^{2}\left(\frac{\sqrt{|B|}a}{v_{0}}\right)}{\sqrt{|B|+\gamma^{2}}-\gamma\,\cosh\left(\frac{2\sqrt{|B|}a}{v_{0}}\right)}\,.$ (112)
@@
$\tilde{a}=\frac{a\gamma}{v_{0}}$ (113)
@@
$\tilde{a}=\frac{a\gamma}{v_{0}}$ (116)
@@
$\lambda(k)=2\tilde{a}^{2}{k}^{2}-\frac{16\tilde{a}^{4}{k}^{3}}{3}+\frac{2}{45}\tilde{a}^{4}\left(368\tilde{a}^{2}+15\right){k}^{4}+O({k}^{5})$ (117)
@@
$\tilde{a}_{c}(\rho)=\sqrt{\frac{3-\sqrt{3}\sqrt{3-\rho^{2}-2\rho}}{4\rho+8}}$ (122)
@@ B.1 Calculation of $I(\rho=1)$
$\lim_{k\to+\infty}k-\lambda(k)=1-\sqrt{1-b}$ (123)
@@
$\lambda=\lambda(b)=\frac{2\left(1-\sqrt{1-b}\right)\sin^{2}(\tilde{a}\sqrt{b})}{\sqrt{1-b}+\cos(2\tilde{a}\sqrt{b})}$ (124)
@@
$\sqrt{1-b_{c}}+\cos(2\tilde{a}\sqrt{b_{c}})=0$ (125)
@@
$\tilde{a}=\frac{1}{2\sqrt{b_{c}}}\arccos(-\sqrt{1-b_{c}})$ (126)
@@
$\lambda(b)\simeq\frac{g(b_{c})}{b_{c}-b}$ (127)
@@
$\simeq 1-\sqrt{1-b_{c}}-\frac{b_{c}-b}{2\sqrt{1-b_{c}}}$
using that $\lambda(b/\lambda^{\prime}(b)\simeq(b_{c}-b)$ From it one obtains
@@
$\tilde{a}=\frac{\cos^{-1}(-1+I(1))}{2\sqrt{I(1)(2-I(1))}}$ (130)
@@
$I(1)=\frac{\pi^{2}}{8\tilde{a}^{2}}\left(1-\frac{1}{\tilde{a}}+O(\frac{1}{\tilde{a}^{2}})\right)=\frac{\pi^{2}D}{4a^{2}}\left(1-\frac{v_{0}^{2}}{\gamma a}+O(\frac{1}{a^{2}})\right)$ (131)
@@
$\lambda=\frac{2(\sqrt{1-b}+1)\sin^{2}(\tilde{a}\sqrt{b})}{\cos(2\tilde{a}\sqrt{b})-\sqrt{1-b}}\quad,\quad k-\lambda=1+\sqrt{1-b}$ (132)
@@
$\cos(2\tilde{a}\sqrt{b_{c}})=\sqrt{1-b_{c}}\quad,\quad\tilde{a}=\frac{1}{2\sqrt{b_{c}}}\arccos(\sqrt{1-b_{c}})$ (133)
@@
$\lambda(b)\simeq\frac{g(b_{c})}{b-b_{c}}$ (134)
@@
$\tilde{a}=\frac{\cos^{-1}(-1+I(1))}{2\sqrt{I(1)(2-I(1))}}$ (136)
@@
$\lambda=\frac{2(\sqrt{1+|b|}+1)\sinh^{2}(\tilde{a}\sqrt{|b|})}{\sqrt{1+|b|}-\cosh(2\tilde{a}\sqrt{|b|})}\quad,\quad k-\lambda=1+\sqrt{1+|b|}$ (137)
@@
$\cosh(2\tilde{a}\sqrt{|b_{c}|})=\sqrt{1+|b_{c}|}\quad,\quad\tilde{a}=\frac{\cosh^{-1}(\sqrt{1+|b_{c}|})}{2\sqrt{|b_{c}|}}$ (138)
@@
$\lambda(b)\simeq\frac{g(|b_{c}|)}{|b_{c}|-|b|}$ (139)
@@
$\tilde{a}=\frac{\cosh^{-1}(-1+I(1))}{2\sqrt{I(1)(I(1)-2}}$ (141)
@@
$\tilde{a}=\frac{\ln(2I(1))}{2I(1)}(1+\frac{1}{I(1)}-\frac{1}{I(1)\ln(2I(1))}+\dots)$ (142)
@@ Appendix C Survival probability in an interval
Let us denote $S_{\pm}(x,t)$ the probability that a RTP initially in the interval $[-a,a]$ remains inside this interval for all times up to time $t$, and is at position $x$ in state $\sigma$ at time $t$. At large $t$ one expects that $S_{\pm}(x,t)\sim S_{\pm}(x)e^{-rt}$. Here $r>0$ is the survival rate $r=-\lim_{T\to+\infty}\ln S(T)$, which does not depend on the details of the initial condition. It is given by the smallest eigenvalue of the linear problem
$-rS_{\sigma}(x)=-v_{0}\sigma\partial_{x}S_{\sigma}(x)+\gamma S_{-\sigma}(x)-\gamma S_{\sigma}(x)$ (144)
with $\sigma=\pm 1$, with boundary conditions $S_{+}(-a)=S_{-}(a)=0$. One looks for a solution such that $S(x)=S_{+}(x)+S_{-}(x)$ is even in $x$ and $R(x)=S_{+}(x)-S_{-}(x)$ is odd in $x$. The boundary condition becomes simply $S(a)=R(a)$. Let us again the dimensionless units (equivalent to setting $\gamma=v_{0}=1$). One finds that for $r<2$ and $x\in[-a,a]$
$S(x)=A\cos(\alpha x)\quad,\quad R(x)=B\sin(\alpha x)$ (145)
@@
$\tan^{2}(\tilde{a}\sqrt{r(2-r)})=\frac{2-r}{r}$ (146)
@@
$\tilde{a}=\frac{\cos^{-1}(-1+r)}{2\sqrt{r(2-r)}}$ (147)
@@
$S(x)=A\cosh(\alpha x)\quad,\quad R(x)=B\sinh(\alpha x)$ (148)
@@
$\tanh^{2}(\tilde{a}\sqrt{r(r-2)})=\frac{r-2}{r}$ (149)
@@
$\tilde{a}=\frac{\cosh^{-1}(-1+r)}{2\sqrt{r(r-2)}}$ (150)
@@
The functions $S(x)$, upon normalization in the interval $[-a,a]$ then gives the probability to find the particle at position $x$, conditioned upon its long time survival. We see that it takes quite different shapes depending on the value of $\tilde{a}$ with a change of behavior at $\tilde{a}=1/2$. For small intervals the particle will be found closer to the boundaries (quite different from the Brownian), while for large intervals it will be found far from the boundaries (as for the Brownian)
@@
$=\exp\left[-T\left(\gamma-\sqrt{\tilde{\rho}^{2}+\gamma^{2}}+\tilde{\rho}\,\sinh^{-1}{\frac{\tilde{\rho}}{\gamma}}\right)\right]$ (153)

### v2

@@ Appendix B Calculation of SCGF for the occupation time statistics
$k-\lambda(k)=\gamma-\sqrt{\gamma^{2}-B}.$ (105)
@@
$\lambda=\frac{2\gamma\left(\gamma-\sqrt{\gamma^{2}-B}\right)\sin^{2}\left(\frac{\sqrt{B}a}{v_{0}}\right)}{\sqrt{\gamma^{2}-B}+\gamma\cos\left(\frac{2\sqrt{B}a}{v_{0}}\right)},$ (107)
@@
$k-\lambda(k)=\gamma+\sqrt{\gamma^{2}-B},$ (108)
@@
$\lambda=\frac{2\gamma\left(\sqrt{\gamma^{2}-B}+\gamma\right)\sin^{2}\left(\frac{\sqrt{B}a}{v_{0}}\right)}{\gamma\cos\left(\frac{2\sqrt{B}a}{v_{0}}\right)-\sqrt{\gamma^{2}-B}}\,,$ (109)
@@
$k-\lambda(k)=\gamma+\sqrt{\gamma^{2}+|B|},$ (111)
@@
$\lambda=\frac{2\,\gamma\left(\sqrt{|B|+\gamma^{2}}+\gamma\right)\sinh^{2}\left(\frac{\sqrt{|B|}a}{v_{0}}\right)}{\sqrt{|B|+\gamma^{2}}-\gamma\,\cosh\left(\frac{2\sqrt{|B|}a}{v_{0}}\right)}\,,$ (112)
@@
$\tilde{a}=\frac{a\gamma}{v_{0}}.$ (113)
@@
$\tilde{a}=\frac{a\gamma}{v_{0}}.$ (116)
@@
$\lambda(k)=2\tilde{a}^{2}{k}^{2}-\frac{16\tilde{a}^{4}{k}^{3}}{3}+\frac{2}{45}\tilde{a}^{4}\left(368\tilde{a}^{2}+15\right){k}^{4}+O({k}^{5}),$ (117)
@@
$\tilde{a}_{c}(\rho)=\sqrt{\frac{3-\sqrt{3}\sqrt{3-\rho^{2}-2\rho}}{4\rho+8}},$ (122)
@@ B.1 Calculation of $I(\rho=1)$
$\lim_{k\to+\infty}k-\lambda(k)=1-\sqrt{1-b},$ (123)
@@
$\lambda=\lambda(b)=\frac{2\left(1-\sqrt{1-b}\right)\sin^{2}(\tilde{a}\sqrt{b})}{\sqrt{1-b}+\cos(2\tilde{a}\sqrt{b})}.$ (124)
@@
$\sqrt{1-b_{c}}+\cos(2\tilde{a}\sqrt{b_{c}})=0.$ (125)
@@
$\tilde{a}=\frac{1}{2\sqrt{b_{c}}}\arccos(-\sqrt{1-b_{c}}).$ (126)
@@
$\lambda(b)\simeq\frac{g(b_{c})}{b_{c}-b},$ (127)
@@
$\simeq 1-\sqrt{1-b_{c}}-\frac{b_{c}-b}{2\sqrt{1-b_{c}}},$
using that $\lambda(b)/\lambda^{\prime}(b)\simeq(b_{c}-b)$ From it one obtains
@@
$\tilde{a}=\frac{\cos^{-1}(-1+I(1))}{2\sqrt{I(1)(2-I(1))}},$ (130)
@@
$I(1)=\frac{\pi^{2}}{8\tilde{a}^{2}}\left(1-\frac{1}{\tilde{a}}+O(\frac{1}{\tilde{a}^{2}})\right)=\frac{\pi^{2}D}{4a^{2}}\left(1-\frac{v_{0}^{2}}{\gamma a}+O(\frac{1}{a^{2}})\right),$ (131)
@@
$\lambda=\frac{2(\sqrt{1-b}+1)\sin^{2}(\tilde{a}\sqrt{b})}{\cos(2\tilde{a}\sqrt{b})-\sqrt{1-b}}\quad,\quad k-\lambda=1+\sqrt{1-b},$ (132)
@@
$\cos(2\tilde{a}\sqrt{b_{c}})=\sqrt{1-b_{c}},\quad\quad\tilde{a}=\frac{1}{2\sqrt{b_{c}}}\arccos(\sqrt{1-b_{c}}).$ (133)
@@
$\lambda(b)\simeq\frac{g(b_{c})}{b-b_{c}},$ (134)
@@
$\tilde{a}=\frac{\cos^{-1}(-1+I(1))}{2\sqrt{I(1)(2-I(1))}}.$ (136)
@@
$\lambda=\frac{2(\sqrt{1+|b|}+1)\sinh^{2}(\tilde{a}\sqrt{|b|})}{\sqrt{1+|b|}-\cosh(2\tilde{a}\sqrt{|b|})},\quad\quad k-\lambda=1+\sqrt{1+|b|},$ (137)
@@
$\cosh(2\tilde{a}\sqrt{|b_{c}|})=\sqrt{1+|b_{c}|},\quad\quad\tilde{a}=\frac{\cosh^{-1}(\sqrt{1+|b_{c}|})}{2\sqrt{|b_{c}|}}.$ (138)
@@
$\lambda(b)\simeq\frac{g(|b_{c}|)}{|b_{c}|-|b|},$ (139)
@@
$\tilde{a}=\frac{\cosh^{-1}(-1+I(1))}{2\sqrt{I(1)(I(1)-2}},$ (141)
@@
$\tilde{a}=\frac{\ln(2I(1))}{2I(1)}(1+\frac{1}{I(1)}-\frac{1}{I(1)\ln(2I(1))}+\dots),$ (142)
@@ Appendix C Survival probability in an interval
Let us denote $S_{\sigma}(x,t)$ the probability that a RTP initially in the interval $[-a,a]$ remains inside this interval for all times up to time $t$, and is at position $x$ in state $\sigma\in\left\{+,-\right\}$ at time $t$. At large $t$ one expects that $S_{\pm}(x,t)\sim S_{\pm}(x)e^{-rt}$. Here $r>0$ is the survival rate $r=-\lim_{T\to+\infty}\ln S(T)$, which does not depend on the details of the initial condition. It is given by the smallest eigenvalue of the linear problem
$-rS_{\sigma}(x)=-v_{0}\sigma\partial_{x}S_{\sigma}(x)+\gamma S_{-\sigma}(x)-\gamma S_{\sigma}(x),$ (144)
with $\sigma=\pm 1$, with boundary conditions $S_{+}(-a)=S_{-}(a)=0$. One looks for a solution such that $S(x)=S_{+}(x)+S_{-}(x)$ is even in $x$ and $R(x)=S_{+}(x)-S_{-}(x)$ is odd in $x$. The boundary condition becomes simply $S(a)=R(a)$. Let us again use the dimensionless units (equivalent to setting $\gamma=v_{0}=1$). One finds that for $r<2$ and $x\in[-a,a]$
$S(x)=A\cos(\alpha x),\quad\quad R(x)=B\sin(\alpha x),$ (145)
@@
$\tan^{2}\left(\tilde{a}\sqrt{r(2-r)}\right)=\frac{2-r}{r},$ (146)
@@
$\tilde{a}=\frac{\cos^{-1}(-1+r)}{2\sqrt{r(2-r)}},$ (147)
@@
$S(x)=A\cosh(\alpha x)\quad,\quad R(x)=B\sinh(\alpha x),$ (148)
@@
$\tanh^{2}\left(\tilde{a}\sqrt{r(r-2)}\right)=\frac{r-2}{r},$ (149)
@@
$\tilde{a}=\frac{\cosh^{-1}(-1+r)}{2\sqrt{r(r-2)}},$ (150)
@@
The function $S(x)$, upon normalization in the interval $[-a,a]$ then gives the probability to find the particle at position $x$ at time $t$, conditioned upon its long time survival up to time $t$. We see that it takes quite different shapes depending on the value of $\tilde{a}$ with a change of behavior at $\tilde{a}=1/2$. For small intervals the particle will be found closer to the boundaries (quite different from the Brownian), while for large intervals it will be found far from the boundaries (as for the Brownian).
@@
$\!\!=\exp\left[-T\left(\gamma-\sqrt{\tilde{\rho}^{2}+\gamma^{2}}+\tilde{\rho}\,\sinh^{-1}{\frac{\tilde{\rho}}{\gamma}}\right)\right],$ (153)
@@
The derivation presented in this appendix is shorter than the derivation of Eq. (58) in the main text. However, the latter derivation presented is straightforward to extend to the case of nonzero drift $\mu\neq 0$, and more generally, the methods that we use in the main text are expected to be applicable in a relatively broad range of scenarios, since they do not rely on exact results at finite $T$.
Appendix E Eigenfunctions for the eigenvalue $\Lambda=0$ in the local time problem
Here we show that, at all $k\geq 0$, $\Lambda=0$ is an eigenvalue of the tilted generator $L_{k}^{\dagger}$ that corresponds to the problem of local time that an RTP spends in the vicinity of the origin. In the spirit of Ref. NT2018DPT in which a very similar situation arises involving drifted Brownian motion, we look for a solution to Eqs. (47) and (48) of the form
$\Phi_{k1}(x)$ $=$ $\begin{cases}A\,e^{\alpha x}+B,&x<0,\\[2.84526pt] 1,&x>0,\end{cases}$ (154)
$\Phi_{k2}(x)$ $=$ $\begin{cases}Ce^{\alpha x},&x<0,\\[2.84526pt] 0,&x>0\end{cases}$ (155)
and with zero eigenvalue. Eqs. (47) and (48) are trivially satisfied at $x>0$, and from the regime $x<0$, they yield the relations
$C$ $=$ $-\mu A/v_{0},$ (156)
$C$ $=$ $-\alpha v_{0}A/\left(2\gamma+\alpha\mu\right),$ (157)
respectively. Comparing the latter two conditions, we obtain $\alpha=2\mu\gamma/\left(v_{0}^{2}-\mu^{2}\right)$. The boundary conditions (49) and (50) at $x=0$ become
$A+B+C$ $=$ $e^{-k}\,,$ (158)
$A+B-C$ $=$ $e^{k}\,.$ (159)
Eqs. (156), (158) and (159) are a set of linear equations for the coefficients $A,B,C$.
The solution to these equations is
$A$ $=$ $\frac{v_{0}}{\mu}\sinh k,$ (160)
$B$ $=$ $\cosh k-\frac{v_{0}}{\mu}\sinh k,$ (161)
$C$ $=$ $-\sinh k\,.$ (162)
We note that $A>0$ at all $k>0$, and $B>0$ at $0<k<k_{c}$ where $k_{c}$ is the criticial point given in Eq. (62) of the main text. We thus find that in the subcritical regime $0<k<k_{c}$, $\Phi_{k1}$ in Eq. (154) is positive. In the supercritical the solution Eq. (154) ceases to be positive for all $x$, but the physical solution in this regime is the localized eigenfunction given by Eq. (51) of the main text.
The solution given in this appendix, with eigenvalue $\Lambda=0$, exists at all $k\geq 0$, and so does the solution given in the main text, with the eigenvalue given in Eq. (56). For nonzero drift, as explained in the main text, the true SCGF $\Lambda(k)$ is given by Eq. (61), which is the largest eigenvalue of the tilted operator.